OpenStudy (anonymous):

find the transition matrix corresponding to the change of basis from {v1,v2} to {u1,u2} where v1= {5,2}^t v2 ={7,3}^t u1=[3,2]^t u2=[1,1]^t these are all vertical btw

OpenStudy (zarkon):

\[\left[\begin{matrix}3 & 1 \\ 2 & 1\end{matrix}\right]^{-1}\left[\begin{matrix}5 & 7 \\ 2 & 3\end{matrix}\right]=\left[\begin{matrix}3& 4 \\ -4 & -5\end{matrix}\right]\]

OpenStudy (anonymous):

here u found inverse of u? 1/det.adj?

OpenStudy (zarkon):

you can do it that way...or you can use elementary row operations \[\left[A|I\right]\rightarrow\left[I|A^{-1}\right]\]

OpenStudy (anonymous):

which way is this one? EROs?

OpenStudy (anonymous):

identity matrix to plot graph? o_O? im lost now

OpenStudy (anonymous):

sorry im looking at both my questions at once i believe i knw how to plot it now

OpenStudy (zarkon):

lol ;)

OpenStudy (anonymous):

now i feel silly ;)

OpenStudy (anonymous):

dont remember how do u find adjoint of a 2x2?

OpenStudy (anonymous):

dnt u transpose for adjoint, change the rows to colomn?